limz→0 z^{2}-sinz/z


1
Expert's answer
2021-08-11T07:28:40-0400

"\lim _{z\to \:0}\left(z^2-\frac{\sin \left(z\right)}{z}\right)\\
\lim _{x\to a}\left[f\left(x\right)\pm g\left(x\right)\right]=\lim _{x\to a}f\left(x\right)\pm \lim _{x\to a}g\left(x\right)\\
With\:the\:exception\:of\:indeterminate\:form\\
=\lim _{z\to \:0}\left(z^2\right)-\lim _{z\to \:0}\left(\frac{\sin \left(z\right)}{z}\right)\\
=0-1\\
=-1"